# Pseudo-labelling meets Label Smoothing for Noisy Partial Label Learning

### v2

@@ 4.2 Results and Discussion
SOTA comparisons: Table 1 compares PALS with the SOTA methods on CIFAR-10 and CIFAR-100 datasets. The results are presented with three different partial rates and two different noise rates, comprising a total of six combinations. PALS surpasses all other methods in all settings, providing clear evidence of its effectiveness. It is also apparent that the PLL approaches struggle to generalize effectively in noisy settings, highlighting the necessity for noise-resistant adaptations.
@@
Fine-grained classification: We first evaluate PALS on the CIFAR-100H dataset, which considers label hierarchy and divides the CIFAR-100 dataset into 20 coarse labels (5 fine-grained labels per coarse label). When curating the NPLL benchmark, the candidate sets are restricted to lie within the same coarse label. PALS obtains the best results in this modified setting (Table 4).
@@
Crowdsourced Datasets: Prior efforts of classification on crowd-sourced datasets [25] take a data-centric approach, involving label enhancement through semi-supervised and noisy label learning, with subsequent evaluation by training a model on the enhanced dataset. Instead, we tackle this as a PLL/NPLL problem and employ PALS. As the only key change in the framework, we utilize the normalized annotation label frequency per image in the weighted KNN step. We construct the partial label by taking into account the annotations provided by human annotators. Table 5(a) presents results under the assumption of ten available human annotators, simulating the PLL scenario with a negligible noise rate, while Table 5(b) displays results assuming three human annotators, simulating the NPLL scenario characterized by a substantial noise rate. We compare against two noisy label learning methods: Divide-Mix [14] and ELR+ [16], a PLL method (PiCO) and an NPLL method (ALIM). In all cases, we use a pre-trained ResNet50 backbone to allow a fair one-to-one comparison. PALS achieves competitive performance on Treeversity. On Benthic and Plankton, which contain underwater images, there is a sharp decline in the performance of PiCO and ALIM. PALS displays promising generalization capability by outperforming them by a significant margin.
@@ 4.3 Ablation studies
In Table 7(e), we perform ablations using two different noise rates: ($\eta=0.3$ and $\eta=0.5$), while setting $q=0.05$. All ablations are performed using the CIFAR-100 dataset.
@@ 5 Conclusion
We propose PALS, a minimalistic framework for NPLL that iteratively performs pseudo-labelling and classifier training. Pseudo-labelling involves the weighted KNN algorithm, and a classifier is trained using the labelled samples while leveraging label smoothing and optional regularization components. PALS achieves SOTA results in a comprehensive set of experiments, varying partial rates and noise rates. The gains are more pronounced with an increase in noise and the number of classes. Notably, PALS excels in fine-grained classification tasks. Unlike many previous methods, PALS preserves its performance as the ambiguity in the dataset increases. PALS also generalizes well to the real world crowd-sourced datasets.

### v3

@@ 4.2 Results and Discussion
SOTA comparisons: Table 1 compares PALS with the SOTA methods on CIFAR-10 and CIFAR-100 datasets. The results are presented with three different partial rates and three different noise rates, comprising a total of nine combinations. PALS surpasses all other methods in all settings, providing clear evidence of its effectiveness. It is also apparent that the PLL approaches struggle to generalize effectively in noisy settings, highlighting the necessity for noise-resistant adaptations.
@@
Simulated fine-grained datasets: We first evaluate PALS on the CIFAR-100H dataset, which considers label hierarchy and divides the CIFAR-100 dataset into 20 coarse labels (5 fine-grained labels per coarse label). When curating the NPLL benchmark, the candidate sets are restricted to lie within the same coarse label. PALS obtains the best results in this modified setting (Table 4).
@@
Fine-grained crowd-sourced datasets: Prior efforts of classification on crowd-sourced datasets [21] take a data-centric approach, involving label enhancement through semi-supervised and noisy label learning, with subsequent evaluation by training a model on the enhanced dataset. Instead, we tackle this as a PLL/NPLL problem and employ PALS. As the only key change in the framework, we utilize the normalized annotation label frequency per image in the weighted KNN step. We construct the partial label by taking into account the annotations provided by human annotators. Table 5(a) presents results under the assumption of ten available human annotators, simulating the PLL scenario with a negligible noise rate, while Table 5(b) displays results assuming three human annotators, simulating the NPLL scenario characterized by a substantial noise rate. We compare against two noisy label learning methods: Divide-Mix [12] and ELR+ [14], a PLL method (PiCO) and an NPLL method (ALIM). In all cases, we use a pre-trained ResNet50 backbone to allow a fair one-to-one comparison. PALS achieves competitive performance on Treeversity. On Benthic and Plankton, which contain underwater images, there is a sharp decline in the performance of PiCO and ALIM. PALS displays promising generalization capability by outperforming them by a significant margin.
@@
Pseudo $q=0.05$ $q=0.05$
label $\eta=0.3$ $\eta=0.5$
76.71 72.86
77.61 75.15
(f) Impact of Eq. 1
Table 7: Ablation studies
@@ 4.3 Ablation studies
In Table 7, we perform ablations using two different noise rates: ($\eta=0.3$ and $\eta=0.5$), while setting $q=0.05$. All ablations are performed using the CIFAR-100 dataset.
@@
Impact of pseudo-label computation in Eq. 1: The first row of Tab. 7(f) shows the results when Eq. 1 is skipped, and class posterior probabilities are computed directly based on the partial label $Y_{k}$. There is a small drop in performance, showing that initial pseudo-label computation is important for robustness against noise.
@@
Training time: PALS is about 1.5 times faster to train compared to PiCO+ and ALIM. On an NVIDIA GeForce RTX 2080 Ti GPU, one epoch of PALS takes 40 seconds, while one epoch of ALIM takes 60 seconds. The numbers reported in our paper are by training for 500 epochs, while ALIM reports numbers after training for 800-1000 epochs (based on their official code). So effectively, PALS is about three times faster to train. On larger models, the gap widens. For example, for a WideResNet, PALS takes 221 seconds, and ALIM takes 550 seconds per epoch, making PALS 2 times faster.
@@ 5 Conclusion
We propose PALS, a framework for NPLL that iteratively performs pseudo-labelling and classifier training. Pseudo-labelling involves the weighted KNN algorithm, and a classifier is trained using the labelled samples while leveraging label smoothing and optional regularization components. PALS achieves SOTA results in a comprehensive set of experiments, varying partial rates and noise rates. The gains are more pronounced with an increase in noise and the number of classes. Notably, PALS excels in the simulated fine-grained classification tasks. Unlike many previous methods, PALS preserves its performance as the ambiguity in the dataset increases. PALS also generalizes well to the real world, fine-grained, crowd-sourced datasets.